# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

### v5

@@ Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023; Dettmers et al. 2024) and Multimodal Large Language Models (MLLMs) (Alayrac et al. 2022; Liu et al. 2024a; Bai et al. 2023; Jiao et al. 2024b, a). Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (Jiao et al. 2024a; Yin et al. 2023b). Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses (Leng et al. 2024), potentially threatening reliability and security in applying them in real-world scenarios (Qian et al. 2024; Jiao et al. 2022). To systematically evaluate hallucinations of ImageLLMs, a thread of benchmarks has been developed (Li et al. 2023b; Fu et al. 2023; Liu et al. 2023b). Specifically, POPE (Li et al. 2023b) investigates the hallucination by mining the frequently occurring objects and their combinations. MME (Fu et al. 2023) and MMBench (Liu et al. 2023b) evaluate hallucinations of the existence, attributes, and relations of objects.
@@
In fact, based on the foundation of ImageLLMs, VideoLLMs have shown significant proficiency in understanding the content of single-frame images (Wang et al. 2024). However, their effectiveness diminishes when tasked with comprehending the event, namely the actions performed by humans or objects spanning the entire video (Du et al. 2024; Zelnik-Manor and Irani 2001; Krishna et al. 2017). This limitation is further proved by our results presented in Fig.6. To further evaluate the event-oriented hallucinations of VideoLLMs, two recent benchmarks (Wang et al. 2024; Guan et al. 2024) develop distinct event comprehension tasks. Specifically, HallusionBench (Guan et al. 2024) reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer (Wang et al. 2024) selects long videos containing multiple events and then queries the model to evaluate their chronological orders. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to probe the model’s authentic ability to comprehend the complex events [^1] A more detailed analysis of event-related VQA samples in HallusionBench (Guan et al. 2024) and VideoHallucer (Wang et al. 2024) is provided in the supplementary materials due to page limitations..
@@
When delving into event-related hallucinations in VideoLLMs, we observe that such errors originate from a consistent issue across different training stages: imbalanced training data distributions, which inject prior knowledge and bias into the model. Initially, building upon LLM backbones, VideoLLMs memorize events with higher frequencies purely through textual statistics, thereby forming strong language-driven priors (Agrawal, Batra, and Parikh 2016; Agarwal, Shetty, and Fritz 2020). For instance, the commonly described event “person drinks coffee” frequently appears in textual corpora, establishing a misleading prior that persists regardless of visual evidence (Fig.1, left). Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, similar imbalances surface during the vision-language alignment stage, where frequently co-occurring visual-textual pairs reinforce spurious object-event associations. As shown in the right column of Fig.1, both Video-LLaVA (Lin et al. 2023) and ShareGPT4Video (Chen et al. 2024) intuitively conclude “people are riding the bicycle” simply from the presence of a person and a bicycle in the scene, disregarding the actual video context.
@@
Video Large Language Models. Large Language Models (LLMs) have achieved remarkable success in traditional NLP tasks and exhibit strong generalization to user-defined instructions, enabling them to act as versatile assistants. Building on this foundation, Multimodal Language Models (MLLMs) extend LLMs with visual inputs, facilitating unified text-vision understanding. To further incorporate video comprehension and temporality, video-supported MLLMs have emerged. Methods such as Video-ChatGPT (Maaz et al. 2023) apply spatial-temporal pooling, while Video-LLaMA (Zhang, Li, and Bing 2023) introduces a Video Q-Former for frame aggregation. Vista-LLaMA (Ma et al. 2023) enhances temporal modeling via a temporal Q-Former projector, and LLaMA-VID (Li, Wang, and Jia 2023) proposes a dual-token paradigm to represent both context and content per frame. Although these approaches perform well on benchmarks, they still struggle with hallucinations in temporal understanding.
Video Evaluation Benchmarks. Despite the rapid progress of MLLMs, a comprehensive evaluation of their capabilities remains essential. For image MLLMs, various benchmarks have been developed to assess perception, reasoning, and cognition (Fu et al. 2023; Liu et al. 2023b; Li et al. 2024c). With the emergence of VideoLLMs, benchmarks like TempCompass (Liu et al. 2024b) and FunQA (Xie et al. 2023) have been proposed to evaluate temporal perception and complex video comprehension.
@@
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs, arising from factors such as imbalanced training data, weak attention to visual content, and other possible reasons. To mitigate hallucinations in image MLLMs, several strategies have been proposed. Woodpecker (Yin et al. 2023a) introduces a post-hoc correction mechanism leveraging additional visual cues, while Volcano (Lee et al. 2023) adopts a critique-revise-decide framework for iterative self-revision. From a data perspective, LRV-Instruction (Liu et al. 2023a) and HalluciDoctor (Yu et al. 2024) aim to alleviate hallucinations by constructing more balanced, high-quality instruction tuning datasets.
Another line of work addresses hallucinations during inference, where contrastive decoding (CD) has shown effectiveness without requiring additional training. VCD (Leng et al. 2024) mitigates statistical bias and language priors by contrasting original and distorted images during decoding. However, existing CD methods mainly target ImageLLMs, with limited exploration in the video domain. Motivated by this gap, we propose Temporal Contrastive Decoding (TCD) to reduce event hallucinations in VideoLLMs.
@@
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g., ActivityNet (Caba Heilbron et al. 2015), (Das et al. 2013), (Soomro 2012)) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
@@
Binary question answering. Following previous benchmarks (Wang et al. 2024; Li et al. 2023b), we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o(Hurst et al. 2024), using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks (Liu et al. 2024b). We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
@@
VideoChatGPT (Maaz et al. 2023) 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 (Li et al. 2024b) 16.67 4.59 47.67 1.55 22.55 32.76 2.64
PLLaVA (Xu et al. 2024) 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID (Li, Wang, and Jia 2023) 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video (Li et al. 2024a) 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA (Lin et al. 2024) 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video (Chen et al. 2024) 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA (Lin et al. 2023) 30.70 8.26 57.51 7.25 41.18 45.97 7.62
@@
GPT-4V (Achiam et al. 2023) 57.89 11.01 79.27 16.58 92.16 76.53 14.57
Gemini-1.5-Pro (Team et al. 2023) 61.40 49.54 83.42 39.90 96.08 80.44 43.38
GPT-4o (Hurst et al. 2024) 83.33 58.72 92.75 41.97 100.00 91.93 48.01
@@
To mitigate the aforementioned hallucinations in VideoLLMs, the most straightforward solution is to enhance finetuning with additional debiased data. However, this approach is highly resource-intensive, as it requires extensive efforts to collect such videos and annotate corresponding QA pairs. Therefore, we take inspiration from (Leng et al. 2024) and propose our Temporal Contrastive Decoding (TCD) method, which alleviates event-oriented hallucinations in VideoLLMs by comparing the logits of the original video and its counterpart during the decoding stage.
@@
Evaluation Setup. We test 8 open-source Video and 3 closed-source VideoLLMs on our EventHallusion. Open-source VideoLLMs comprise VideoChatGPT (Maaz et al. 2023), VideoChat2 (Li et al. 2024b), PLLaVA (Xu et al. 2024), LLaMA-VID (Li, Wang, and Jia 2023), LLaVA-NeXT-Video (Li et al. 2024a), VILA (Lin et al. 2024), ShareGPT4Video (Chen et al. 2024), and Video-LLaVA (Lin et al. 2023). For closed-source VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (Achiam et al. 2023), GPT-4o(Hurst et al. 2024), and Gemini-1.5-Pro (Team et al. 2023), to conduct the systematic comparison of the robustness between open-source and closed-source MLLMs.
@@
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (Li et al. 2024a), VideoChat2 (Li et al. 2024b), and VILA (Lin et al. 2024) as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA (Liu et al. 2024a) based on the projection layer design, while VideoChat2 adopts the Q-Former (Li et al. 2023a) to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
@@
We apply our proposed TCD method on three prevalent VideoLLMs, namely LLaVA-NeXT (Li et al. 2024a), VideoChat2 (Li et al. 2024b), and VILA (Lin et al. 2024), and evaluate their performances on EventHallusion. As shown in Tab.2, equipping with TCD, three VideoLLMs achieve a consistent performance boost. This demonstrates that TCD can be seamlessly integrated with versatile VideoLLMs and reduce their event-related hallucinations without bells and whistles.
@@
We further explore the effects of the temporal downsampling ratio when constructing the contrastive video in our TCD. Specifically, we adopt LLaVA-NeXT-Video(Li et al. 2024a) as the backbone model with 32 frames sampled for the original video input and downsample the number of sampled frames to construct the contrastive video. This experiment is conducted on the binary classification task of EventHallusion. From Tab.4, we have the following observations: (1) Extreme temporal downsampling (i.e., 1 frame) can lead to severe performance degradation of our TCD approach. This phenomenon may arise because single-frame input significantly diminishes spatial cues, leading to distorted prediction logits during contrastive learning. (2) A low temporal downsampling rate (e.g., 16 frames) performs worse than higher rates (e.g., 4 and 8 frames). This indicates that narrowing the gap between the contrastive video and the original one weakens the advantages of our TCD approach.

### v6

@@ Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (brown2020language; touvron2023llama; dettmers2024qlora) and Multimodal Large Language Models (MLLMs) (alayrac2022flamingo; liu2024visual; bai2023qwen; jiao2024lumen; jiao2024rode). Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (jiao2024rode; yin2023foodlmm). Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses (leng2024vcd), potentially threatening reliability and security in applying them in real-world scenarios (qian2024nuscenes; jiao2022more). To systematically evaluate hallucinations of ImageLLMs, a thread of benchmarks has been developed (pope; fu2023mme; liu2023mmbench). Specifically, POPE (pope) investigates the hallucination by mining the frequently occurring objects and their combinations. MME (fu2023mme) and MMBench (liu2023mmbench) evaluate hallucinations of the existence, attributes, and relations of objects.
@@
In fact, based on the foundation of ImageLLMs, VideoLLMs have shown significant proficiency in understanding the content of single-frame images (wang2024videohallucer). However, their effectiveness diminishes when tasked with comprehending the event, namely the actions performed by humans or objects spanning the entire video (du2024towards; zelnik2001event; krishna2017dense). This limitation is further proved by our results presented in Fig.6. To further evaluate the event-oriented hallucinations of VideoLLMs, two recent benchmarks (wang2024videohallucer; guan2024hallusionbench) develop distinct event comprehension tasks. Specifically, HallusionBench (guan2024hallusionbench) reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer (wang2024videohallucer) selects long videos containing multiple events and then queries the model to evaluate their chronological orders. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to probe the model’s authentic ability to comprehend the complex events [^1] A more detailed analysis of event-related VQA samples in HallusionBench (guan2024hallusionbench) and VideoHallucer (wang2024videohallucer) is provided in the supplementary materials due to page limitations..
@@
When delving into event-related hallucinations in VideoLLMs, we observe that such errors originate from a consistent issue across different training stages: imbalanced training data distributions, which inject prior knowledge and bias into the model. Initially, building upon LLM backbones, VideoLLMs memorize events with higher frequencies purely through textual statistics, thereby forming strong language-driven priors (agrawal2016analyzing; agarwal2020towards). For instance, the commonly described event “person drinks coffee” frequently appears in textual corpora, establishing a misleading prior that persists regardless of visual evidence (Fig.1, left). Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, similar imbalances surface during the vision-language alignment stage, where frequently co-occurring visual-textual pairs reinforce spurious object-event associations. As shown in the right column of Fig.1, both Video-LLaVA (videollava) and ShareGPT4Video (chen2024sharegpt4video) intuitively conclude “people are riding the bicycle” simply from the presence of a person and a bicycle in the scene, disregarding the actual video context.
@@
Video Large Language Models. Large Language Models (LLMs) have achieved remarkable success in traditional NLP tasks and exhibit strong generalization to user-defined instructions, enabling them to act as versatile assistants. Building on this foundation, Multimodal Language Models (MLLMs) extend LLMs with visual inputs, facilitating unified text-vision understanding. To further incorporate video comprehension and temporality, video-supported MLLMs have emerged. Methods such as Video-ChatGPT (maaz2023videochatgpt) apply spatial-temporal pooling, while Video-LLaMA (zhang2023videollama) introduces a Video Q-Former for frame aggregation. Vista-LLaMA (ma2023vista) enhances temporal modeling via a temporal Q-Former projector, and LLaMA-VID (llamavid) proposes a dual-token paradigm to represent both context and content per frame. Although these approaches perform well on benchmarks, they still struggle with hallucinations in temporal understanding.
Video Evaluation Benchmarks. Despite the rapid progress of MLLMs, a comprehensive evaluation of their capabilities remains essential. For image MLLMs, various benchmarks have been developed to assess perception, reasoning, and cognition (fu2023mme; liu2023mmbench; li2024eyes). With the emergence of VideoLLMs, benchmarks like TempCompass (liu2024tempcompass) and FunQA (xie2023funqa) have been proposed to evaluate temporal perception and complex video comprehension.
@@
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs, arising from factors such as imbalanced training data, weak attention to visual content, and other possible reasons. To mitigate hallucinations in image MLLMs, several strategies have been proposed. Woodpecker (yin2023woodpecker) introduces a post-hoc correction mechanism leveraging additional visual cues, while Volcano (lee2023volcano) adopts a critique-revise-decide framework for iterative self-revision. From a data perspective, LRV-Instruction (lrv_instruction) and HalluciDoctor (yu2024hallucidoctor) aim to alleviate hallucinations by constructing more balanced, high-quality instruction tuning datasets.
Another line of work addresses hallucinations during inference, where contrastive decoding (CD) has shown effectiveness without requiring additional training. VCD (leng2024vcd) mitigates statistical bias and language priors by contrasting original and distorted images during decoding. However, existing CD methods mainly target ImageLLMs, with limited exploration in the video domain. Motivated by this gap, we propose Temporal Contrastive Decoding (TCD) to reduce event hallucinations in VideoLLMs.
@@
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g., ActivityNet (caba2015activitynet), (das2013thousand), (soomro2012ucf101)) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
@@
Binary question answering. Following previous benchmarks (wang2024videohallucer; pope), we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o(hurst2024gpt4o), using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks (liu2024tempcompass). We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
@@
VideoChatGPT (maaz2023videochatgpt) 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 (li2024mvbench) 16.67 4.59 47.67 1.55 22.55 32.76 2.64
PLLaVA (xu2024pllava) 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID (llamavid) 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video (li2024llava) 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA (lin2024vila) 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video (chen2024sharegpt4video) 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA (videollava) 30.70 8.26 57.51 7.25 41.18 45.97 7.62
@@
GPT-4V (achiam2023gpt) 57.89 11.01 79.27 16.58 92.16 76.53 14.57
Gemini-1.5-Pro (team2023gemini) 61.40 49.54 83.42 39.90 96.08 80.44 43.38
GPT-4o (hurst2024gpt4o) 83.33 58.72 92.75 41.97 100.00 91.93 48.01
@@
To mitigate the aforementioned hallucinations in VideoLLMs, the most straightforward solution is to enhance finetuning with additional debiased data. However, this approach is highly resource-intensive, as it requires extensive efforts to collect such videos and annotate corresponding QA pairs. Therefore, we take inspiration from (leng2024vcd) and propose our Temporal Contrastive Decoding (TCD) method, which alleviates event-oriented hallucinations in VideoLLMs by comparing the logits of the original video and its counterpart during the decoding stage.
@@
Evaluation Setup. We test 8 open-source Video and 3 closed-source VideoLLMs on our EventHallusion. Open-source VideoLLMs comprise VideoChatGPT (maaz2023videochatgpt), VideoChat2 (li2024mvbench), PLLaVA (xu2024pllava), LLaMA-VID (llamavid), LLaVA-NeXT-Video (li2024llava), VILA (lin2024vila), ShareGPT4Video (chen2024sharegpt4video), and Video-LLaVA (videollava). For closed-source VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (achiam2023gpt), GPT-4o(hurst2024gpt4o), and Gemini-1.5-Pro (team2023gemini), to conduct the systematic comparison of the robustness between open-source and closed-source MLLMs.
@@
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (li2024llava), VideoChat2 (li2024mvbench), and VILA (lin2024vila) as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA (liu2024visual) based on the projection layer design, while VideoChat2 adopts the Q-Former (li2023blip) to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
@@
We apply our proposed TCD method on three prevalent VideoLLMs, namely LLaVA-NeXT (li2024llava), VideoChat2 (li2024mvbench), and VILA (lin2024vila), and evaluate their performances on EventHallusion. As shown in Tab.2, equipping with TCD, three VideoLLMs achieve a consistent performance boost. This demonstrates that TCD can be seamlessly integrated with versatile VideoLLMs and reduce their event-related hallucinations without bells and whistles.
@@
We further explore the effects of the temporal downsampling ratio when constructing the contrastive video in our TCD. Specifically, we adopt LLaVA-NeXT-Video(li2024llava) as the backbone model with 32 frames sampled for the original video input and downsample the number of sampled frames to construct the contrastive video. This experiment is conducted on the binary classification task of EventHallusion. From Tab.4, we have the following observations: (1) Extreme temporal downsampling (i.e., 1 frame) can lead to severe performance degradation of our TCD approach. This phenomenon may arise because single-frame input significantly diminishes spatial cues, leading to distorted prediction logits during contrastive learning. (2) A low temporal downsampling rate (e.g., 16 frames) performs worse than higher rates (e.g., 4 and 8 frames). This indicates that narrowing the gap between the contrastive video and the original one weakens the advantages of our TCD approach.